OpenStudy (anonymous):

PLEASE HELP Find an explicit rule for the nth term of the sequence. -5, -25, -125, -625, ...

OpenStudy (anonymous):

an = 5 • -5n an = 5 • -5n + 1 an = -5 • 5n an = -5 • 5n - 1

OpenStudy (welshfella):

first term = -5 second term = -5 * 5 third term = -5 * 5^2 4th term = -5 * 5^3 can you see the pattern?

OpenStudy (anonymous):

Yes as the terms increase so does the exponent

OpenStudy (welshfella):

yes and if the term is 2 the exponent is 1 and if term is 3 the exponent is 2 for the first term the exponent is 0

OpenStudy (anonymous):

Would it be C??

OpenStudy (anonymous):

@welshfella

OpenStudy (welshfella):

no because the n = exponent

OpenStudy (anonymous):

That's the consistent pattern that's occurring, right

OpenStudy (anonymous):

Oh

OpenStudy (anonymous):

Ohhhh wait

OpenStudy (welshfella):

its has to be c or d because of the -5

OpenStudy (anonymous):

Can't it be B because of the continuous addition to the exponent as the terms increase?

OpenStudy (welshfella):

no because the first term is -5 not 5

OpenStudy (anonymous):

Then it must be C because the exponent is being subtracted from, correct?

OpenStudy (anonymous):

*isn't

OpenStudy (welshfella):

the way you have written the options is confusing they should be 5 * (-5)^n 5 * (-5)^(n+1) -5 * (5)^n -5 * (5)^(n-1)

OpenStudy (welshfella):

check out a n = 1:- a 1 = first term = 5*(-5)^1 = -25

OpenStudy (welshfella):

try d when n = 1

OpenStudy (welshfella):

a1 = -5 * (5)^(1-1) = -5 ^ 5^0 = -5 * 1 = -5

OpenStudy (welshfella):

* that should be -5 * 5^0

OpenStudy (welshfella):

do you follow that ok?

OpenStudy (anonymous):

Did you use D in your last problem?

OpenStudy (welshfella):

a2 = -5*(5)^(2-1) = -5 * 5^1 = -5 * 5 = -25

OpenStudy (welshfella):

yes

OpenStudy (anonymous):

Would it be D then since the answer is -5?????

OpenStudy (welshfella):

i used d for both calculations

OpenStudy (anonymous):

*?

OpenStudy (anonymous):

Oh I just saw it

OpenStudy (anonymous):

Thank you so much

OpenStudy (welshfella):

well you have to check a few not just the first as you see it works for n = 1 and n = 2

OpenStudy (anonymous):

Would you mind helping me with 2 more

OpenStudy (welshfella):

normally i would but i gotta go right now

OpenStudy (anonymous):

Oh alright. Thanks again!